OpenStudy (anonymous):

Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points)

OpenStudy (anonymous):

solution: x = -2 y = -4 --- part A: --- 5x-5y=10 (3x-2y=2) * 2 --- 5x-5y=10 6x-4y=4 --- add: 11x-9y=14 --- new system: 11x-9y=14 3x-2y=2